OpenStudy (yellowlegoguy99):

A function has the rule "multiply by two, and then subtract three." What is the output when the input is 8? ten thirteen eighteen twenty-two

OpenStudy (yellowlegoguy99):

Medal to who answers best? :D

OpenStudy (yellowlegoguy99):

I have more questions :/

OpenStudy (anonymous):

f(x)=2x-3 So, f(8) = 2*8-3 = 13

OpenStudy (yellowlegoguy99):

Kay ty

OpenStudy (yellowlegoguy99):

Mickey bought 2 books from the bargain bin at his school's book fair and paid four dollars. If buying books from the bargain bin follows a direct variation, how much did his sister pay for the 6 books she bought? eight dollars twelve dollars twenty-four dollars thirty-six dollars

OpenStudy (anonymous):

next question in another post please